OpenStudy (anonymous):

PLEASE HELP ME!!! Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). 1. y = 2x + 6 2. y = -1/2x + 3 3. y = -1/2x + 6 4. y = 2x + 3

OpenStudy (mrnood):

you line y= 2x+2 is in the standard form of a straight line y=mx + b in the standard equation m is the gradient so FIRST tell me what is the gradient of your line y = 2 x + 2 y = m x + b (m= gradient)

OpenStudy (anonymous):

So, 2 is the gradient

OpenStudy (mrnood):

Yes well done 2 is the gradient OK - I think that there is a mistake in your question, or how you posted it none of the answers is correct. Can you check what you posted please?

OpenStudy (anonymous):

Yes, there was a problem in the way I posted it. I just fixed the answers, so hopefully they're correct now

OpenStudy (mrnood):

a line which is perpendicular to the line y= mx + b has a gradient -1/m in you case this would be -1/2 so the equation would be \[y =\frac{ -1 }{ 2 }x +b\]

OpenStudy (mrnood):

OK good so now you can see that 2 of your answers have a gradient of -1/2

OpenStudy (anonymous):

So the answer would be y = -1/2x + 2?

OpenStudy (mrnood):

no

OpenStudy (anonymous):

Oh, so how do I get the correct value for b?

OpenStudy (mrnood):

you know the equation is y = -1/2 x+ b and you know that the point x=6, y=3 is on the line SO 3 = -1/2 (6) + b so work out what b is

OpenStudy (anonymous):

Okay, I solved for b and ended up with 6. So the equation would be y = -1/2x + 6, which is #3. Correct?

OpenStudy (mrnood):

well done - yes

OpenStudy (anonymous):

Thanks so much, that was amazing help. I'm going to fan you and give you a medal :)

OpenStudy (mrnood):